OpenStudy (doodleduck8):

Solve the equation for y: 4x + 3y = -12

OpenStudy (osprey):

Can't SOLVE it, but CAN rearrange it ... 4x + 3y = -12 3y=-4x-12 y=-4/3 x -4 To "solve" you'd need a value of x to solve for y, or a value of y to solve for x ... http://perendis.webs.com

OpenStudy (kellygirl):

why do you have a math question on an English study thing